North Eastern Locomotive Preservation Group

  • Increase font size
  • Default font size
  • Decrease font size
Q6 Blog

A record of events, activities and our work with the Q6



Wednesday 21st June 2017

Winter Maintanance reaches Mid-Summer

A warm muggy day at Grosmont today, with Derek Shorten, Jon Bradley, Nigel Hall & Ian Pearson. Also Gordon Wells fitting conduit and a couple of sockets in the pit.

After a cup of tea a slow start was made on progressing today's work load, Derek Setting up the lighting down in the running shed and taking the equipment needed for final cleaning and painting of the L/H frames. Jon & I replaced the remaining Klinger packing top R/H. Then onto boiler dome to remove the broken stud. This has had a nut welded onto it, so eventually we found a socket that fitted this nut 30mm . A long pipe was attached to the socket bar and after about 4 attempts to remove the stud it sheared again. I have left this job until Saturday as it may need drilling out ,plus the fact there was someone working in the boiler assisting the boiler smith tapping the flue holes in the tube plate. In the meantime Nigel got started with the painting and by the end of the day the left hand side is now all primed.

Lunch was going to have been on the bench outside Deviation shed but as Nigel was about to take a bite of his sandwich the heaven's opened torrential rain came down so lunch was inside Deviation shed. The rain was hammering down and it wasn't long before it was flowing under the side door through the workshop and out into the pit.

After lunch the remaining four Spring hanger adjuster bolts which Chris Parrish was re- cutting the damaged threads were brought back to the workshop and are now on the bench with the others, he still has one nut that needs some attention but will be ready on Saturday. Had a word with Piglet re NDT of the hanger bolts he is trying to get this done soon!

At long last the L/H back plate has been removed ,thanks to Shaun Bowler and his mate (railway fitters). it was taken to the press with about 10 tons of pressure and quite a bit of heat it eventually started to move free. No wonder it would not move the amount of rust holding it. I don't think it had been off for many years! see photo's, Jon and I started to clean the back plate and spindle still needs some more cleaning.

The valve spindle after dismantling showing keyway - Ian Pearson

The backplate showing keyway - Ian Pearson

We had a look in the White metal shop where our axle boxes are being white metaled about six are done.

One of the newly metalled axle boxes - Ian Pearson

 

Saturday 17th June 2017

Extremely hot sunny day at Grosmont with Colin Foxton , Peter Ellis& Ian Pearson.

RHS valve plate showing keyway - Ian Pearson

The R/H valve was put in the vice and after half an hour the key was removed, Colin realised that the key had been put in the wrong way. These keys have a slight taper and need to go in taper first this was not so and causing the to jam. These keys seem to be short in depth so they do not hold anything! See photo. When R/H valve was stripped down it was found that the back valve ring has a broken lug. See Photo ,New ring required. The valve has been reassembled and is on the bench for further investigation and repair.

The ring showing the lug broken from 5 o'clock on bench surface - Ian Pearson

The L/H valve has had part of the broken key drilled out by Colin and some heat treatment from Peter and an attempt to remove the back plate from the valve spindle but it wont budge. We may try the press on Wednesday?

Bill Dobson called in for chat and a cup of tea showed him the valves and explained the problems we are having.

I replaced 2 more klinger packing's on the gauge frames ,one left to change top R/H.

Lots of competitors for the railways" race the train" today which started at 2pm on the old track bed opposite Deviation shed , a bit hot for running to Goathland and back to Gosmont! We packed at 2.30 and called it a day.

 

Wednesday 14th June 2017

A hot sunny, but frustrating day at Grosmont with Derek Shorten, Nigel Hall, and myself (eventually, after a call in at Pickering on the way), with Ian Pearson in charge. Bill Dobson joined us for a few hours and Terry Newman paid us a visit at teatime.

Derek and Nigel worked all day in the running shed, Derek cleaning the brackets of the LHS frames and Nigel sanding down both sides of the frames - these are now almost ready for painting. I was found employment filing a small piece of steel to make a new key for the left hand piston valve: still a bit more to do to finish that off - it is on the bench by the RH piston valve. The shed staff have welded a nut onto the broken stud on the dome and Ian had a go at its removal. It would not shift though, and against the possibility of it shearing if more pressure was applied, it was abandoned for the moment. Barney, the chargehand fitter, checked the gauge frame gland nuts. These were assessed as now OK after cleaning the top left hand one with a thread file, and the Klinger packings can be fitted on Saturday.

We had our lunch outside Deviation Shed on the bench, sat in the lovely warm sunshine watching the trains go by and giving Jon Bradley a wave.

Quite a lot of visitors round the shed today: must have been the warm weather!

After lunch Ian and I, with help from Bill, had a look at the left hand valve spindle with the broken key jammed in the key way, but decided to leave it for the time being . So we started to strip the right hand valve to try and get the valve ring free when the nut is tightened by putting extra shims in. After removing the nut and front end of the valve, the cone shaped piece has a key in it, but in spite of several hours effort trying to remove it, it proved impossible to get it out and we had to retire, frustrated and hot.

The NYMR is in the process of making a couple of valve stands, and the conduit for the electric cables for the pit lighting has been delivered and is on the Shed floor by the workshop. Mark O'Brien has yet to finish the firebox flue tube thread recutting, as absolute priority is being given to getting Repton finished (the boiler is now full of water), but his plan is to have that work finished and the completed flue tubes at Grosmont by the time the wheel sets come back from South Devon Engineering.

 

Saturday 10th June 2017

Colin Foxton, Paul Swainston, Nigel Hall, Brian Orange and the JVs.

On arrival the left valve rod was spotted forlornly leaning up against the wall next to the press in the running shed with the last piece of the head and broken key still firmly in place. So back to deviation shed. The rest of the elements of the left valve rod had their mating faces thoroughly cleaned with help from Paul Swainston at the end of his NYMR cleaning shift. A trial assembly of all the elements on to the rod was tried and a similar problem experienced to the right hand rod where once the nut was fully tightened both valve rings were solid. So whether or not the elusive key is removed this issue of the rings not being free on both valves will have to be examined further. Time to go home after varying attempts with penetrating oil, paraffin and gentle persuasion, you guessed it, no movement so the rod was left in the vice outside the workshop with the outside of the head uppermost and the well full of paraffin. Hopefully by next Wednesday or Saturday the situation may have eased.

The valve with the broken key soaking in a pool of paraffin. See Saturday 3rd  - Colin Foxton

In the running shed we continued with the cleaning of the RHS frames, assisted from time to time by Bryan and a succession of JVs. By the end of the day we had them free enough of grease to make the first pass with the orbital sander. Hopefully the LHS is in a fit condition for the orbital sander on Wednesday too.

The rain lashed down. All that was needed was the swarm of midges for it to be an August day in Ft William.

Neal Woods (driving) and Bill Dobson visited us during the day.

 

 

Wednesday 7th June 2017

Todays working party Ian Pearson, Jon Bradley, Derek Shorten & Nigel Hall.

Mark o' Brian asked me if we could re-fit the fire grate to give him better working base in the firebox for the work involved in re- tubing the boiler. Jon & I carried out this task the front section of the grate was moved in a barrow and fitted , we got the Telehandler to bring the back section down to the shed much easier, and got that fitted after lunch. We then removed the oil pipes and brackets from the frames which feed the driving axle boxes these put in store up stairs. Derek & Nigel continued cleaning the right hand frames, still some more cleaning to do. The left hand valve is still in the press no more work has been done on this. Some of the gauge frame gland nuts are quite sloppy on the threads these need to be examined. Four new Klinger Packing's were obtained from the stores last Saturday only one Klinger packing was fitted to the top left gauge frame last week, the rest are on the work bench in deviation shed. No point in fitting any more until the gauge frame gland nuts are sorted.

I will not be at Grosmont this Saturday10/06/17, Nigel intends to go and continue cleaning & painting so contact Nigel Hall first if any one would like a day out at Grosmont shed .

 

 

Tuesday 6th June 2017

D Day!

The South Devon Railway continue to make progress with the wheels. These photographs are of the first wheelset in the lathe for cleaning up the wheel centre. unfortunately there seems to be a lot to clean up.

The first Q6 wheelset in the lathe at Buckfastleigh - Rob Le Chevalier

The first Q6 wheelset in the lathe at Buckfastleigh - Rob Le Chevalier

The first Q6 wheelset in the lathe at Buckfastleigh - Rob Le Chevalier

The first Q6 wheelset in the lathe at Buckfastleigh - Rob Le Chevalier

 

Monday 5th June 2017

The Q6 wheelsets at the South Devon Railway have had their old tyres removed. See SDR Railway Wheels page for background concerning railway wheel assembly.

Q6 whelset at the SDR with the tyres removed and the Gibson rings shown - Rob Le Chevalier

A Q6 wheel set with tyres removed - Rob Le Chevalier

8 off scrap tyres - Rob Le Chevalier

 

Saturday 3rd June 2017

First I would like to mention last Wednesdays report where the steam reverser requires new rings. I was quite shocked to find that after 6 months working on the loco all of a sudden extra work has been found. The glands have already been tightened & lock nuts fitted to the followers, so I marked it off as complete months ago. Good news though, the shed mechanical foreman has volunteered to remove the piston ring & have a new one made. As he had the reverser in bits last year he knows what is required so I took up his offer to do the job.

So to today’s work party Colin Foxton, Ian Pearson, Peter Ellis, Andy & Joan Lowes were in attendance. Nigel Hall sent his apologies as granddad duties were required elsewhere. Mark O’Brian under contract has removed the stub ends to the flue tube bottle ends & we installed the guide into the front tube plate to tap the fire box tube plate holes tomorrow.

Joan Lowes washing up the tea cups after lunch - Andy Lowes

The driver axle boxes have had the old ferrules removed for the oil feed. The new pipes with bushes attached have been put with the boxes ready for when the boxes go back onto the axles. The trailing axle lubrication is staying as it is due to the ash pan; there is very little room to fit anything different to what is there already.

A laminated copy of our Armstrong oiler pads for all NELPG locos has been stuck to the side of the cupboard above the workshop where the stock of pads are kept, just for reference.

The right piston valve has been stripped again to try & get the rings to slide up & down. We have moved 2 shims (.016”) from the back to the front to try & achieve this. It worked until the nut was tightened up fully so the cotter would go into the hole. Both valve rings nipped up solid so will have to be looked at again by someone with more knowledge of this type of valve.

So to the left valve, this is where our problems really did start! When stripping the valve down everything has been laid out on the bench in the order that it came off (please do not move anything). The last piece of the head has a piece of broken key in it that looks as if it has twisted!

The twisted key in the left valve - Andy Lowes

We tried giving it a tap with a copper mallet, no movement. After soaking the joint in penetrating oil & digging out more carbon we tried again with the mallet, you guessed it no movement! This called for drastic action, with the aid of a tape measure we worked out it would just fit into the press in the running shed. (See photo 4894) So the valve rod was set up in the press & 3 then 4 then 5 ton of force was applied I stopped at 10 ton, you are ahead of me now yes no movement!! I then had a word with our friendly MPD manager who was duty fitter this afternoon. He recommended using between 3 & 4 ton of force with a little heat, yet again no movement. The rod was still warm when it was time to leave for home so we could not repeat the procedure again. Having told Paul Middleton what we had done, he said if he had time, he would give it another go later. I hope he has more luck than we did.

Colin Foxton alignes the valve in the running shed press - Andy Lowes

Also today the right side outer frames have been cleaned, well about half were. The grime that has stuck to them is really difficult to remove, to be continued on Wednesday.

Joan Lowes cleaning the RHS frames - Andy Lowes

The gauge frames have been stripped, cleaned & roded through, new packing & stainless steel balls will be fitted to them after the threads have been examined for wear.

I discussed with Paul Middleton the problems we are having with the spring hanger adjusters; I asked him if he could get them NDT inspected, he said he knows someone who comes over a couple of times a week. I asked him to get them checked for us and also all the horn stays as last time the driving wheels were out one was cracked. I also asked about the springs, Paul told me these will be going away to be repaired soon.

The springs - Andy Lowes

Next Wednesday 7-6-17 Ian Pearson will be RO as usual but he is not available next Saturday 10-6-17.This will be my last winter maintenance report, as it is now June it must be summer by now! So as Paul Hutchinson Q6 caretaker called in two weeks ago to Deviation shed to find out where we were up to. I leave the project at a turning point where the loco is stripped & some items refurbished ready for re-assembly. Watch this space for what is happening next Saturday.

It just leaves me to thank everyone who has been involved with the Q6 winter maintenance in whatever capacity whether a one off or most weeks I appreciate your efforts & I hope you enjoyed the experience.

One last thing, Joan & I will try & get over to Deviation shed to work on the Q6 when we can. The next Saturday we have free is 24-6-17, see you all there for mid summer madness!

 

Wednesday 31st May 2017

Another fine sunny day at Grosmont yesterday, although not as warm as last Saturday which made for a better temperature for working. Only Derek Shorten and Ian at the start of the day, although Nigel joined later on, having been stuck in traffic jams, one involving lots of blue lights in Middlesbrough and one at Scaling Dam where they are resurfacing the road with those"dreaded chippings", and I called in to join them for lunch on the bench outside Deviation Shed between other business in the area.

After the usual first cup of tea, Derek set about cleaning the inside of the cab, especially both bottom sides where the wooden floor sits. Ian cleaned the blower valve on the bead blaster and got Owain to skim the valve head. After lunch, Nigel started on the frames preparing them for painting, while Ian cleaned the blow down valve, again with the bead blaster.  Both valves are on the bench, with the  blow down valve requiring a new gasket. Mark O'Brien advised that the NYMR was using copper rings for this, but they did not have one our size in the stores although one can be ordered if they are given the dimensions. Barney also asked if we were going to remove the reverser, or whether he was to do it, as the broad gauge piston ring at the back needed replacing.

There were also rumours floating around that the K1 had failed in Scotland earlier in the week. Checks with the team at Fort William have now confirmed that there are no problems and the K1 is fine.

Working party as usual on Saturday when Paul the weather man is forecasting sunshine but fresher temperatures, so ideal weather for working on the Q6 and watching the trains go by.

 

Sarurday 27th May 2017

On an extremely hot sunny day Ian Pearson, Peter Ellis, Colin Foxton, Joan & Andy Lowes were at Deviation shed working on the Q6. Trevor Wilford called in to catch up with the banter before he had a trip to Pickering on the train so we could wave at him! Gordon Wells called in to progress the pit lighting but requires some conduit now, he did get one of the lights above the workshop working so we have a little more light now up there.

Ian enquired at Armstrong oilers about the pads we use on our locos & was given a list of all the numbers on the pads; I will send a copy to all the loco caretakers etc. for reference.

A number of 110v lamps were repaired & or new tubes/bulbs fitted, all now work.

A spare spring hanger adjuster was found in a box above the workshop, thanks to Nigel Hall who sent an email through the week with a photo showing this.

The end caps were fitted to the injectors; the blower valve & blow down valves have both been removed for refurbishment on the bench.

The axle box pipe work is progressing with the trailing copper lubrication pipes being annealed & the new flexible pipes fitted to the bushes ready for the driving axle.

The tender tank sides & back have been cleaned with kerosene & a survey has been completed on the loco brake gear.

Wednesday 31-5-17

  • Continue with back head fitting refurbishment
  • Both gauge frames to strip, clean & build back up
  • Clean loco & prepare frames for painting

Next weekend Mark O’Brian will be working on the boiler tapping the fire box flue tube holes. This will enable him to measure the threads & order the flue tube bottle ends that are welded onto the tubes. There is about three week lead in time from placing the order to delivery. He uses a supplier who does all the work & includes all the certification job done; all he then has to do is fit them into our boiler.

 
  • «
  •  Start 
  •  Prev 
  •  1 
  •  2 
  •  3 
  •  4 
  •  5 
  •  6 
  •  7 
  •  8 
  •  Next 
  •  End 
  • »


Page 1 of 8